OpenStudy (anonymous):

Help please! Phyllis invested $9,000, a portion earning a simple interest rate of 8 1 4 % per year and the rest earning a rate of 8% per year. After 1 year the total interest earned on these investments was $737.5. How much money did she invest at each rate?

OpenStudy (anonymous):

rate of 8 1/4%

OpenStudy (anonymous):

Im closing this but I still need the help. THank you!

OpenStudy (anonymous):

total amount of money = 9,000 portion at simple interest rate of 8.25 % = x portion at simple interest rate of 8% = 9,000 - x After 1 year: x* 8.25% + (9,000 - x)*8% = 737.5 x * (8.25/100) + (9,000 - x) * 8/100 = 737.5 8.25 x + (9,000 - x) * 8 = 73750 8.25 x + 72,000 - 8 x = 73750 0.25 x = 73750 - 72000 0.25 x = 1,750 x = 7,000

OpenStudy (anonymous):

so 7000 for 8.25 and 2000 for 8%?

OpenStudy (anonymous):

Yes.

OpenStudy (anonymous):

Thats what I got too but its incorrect

OpenStudy (anonymous):

Really? Why? What's the correct solution?

OpenStudy (anonymous):

Are you sure you didn't just forgot a number in the enunciate?

OpenStudy (anonymous):

I got it thanks!!

OpenStudy (anonymous):

You're welcome. :)